A plane takes off from an airport which is at sea level. The aircraft’s altitude in metres at a time of t minutes is modelled by the equation

h = 5400 ln(t + 1)


a) What is the height of the plane after five minutes?

b) What is the rate of climb of the plane after five minutes?

c) Using differentiation, show the curve y = h(t) has no stationary points, and hence comment on how realistic the model would be for large t.


1
Expert's answer
2021-06-08T14:47:15-0400

a)


"h(5)=5400\ln(5+1)=5400\ln(6)\approx9675.5 (m)"

b)


"h'(t)=\dfrac{5400}{t+1}"

"h'(5)=\dfrac{5400}{5+1}=900 (m/min)"

c)

"h(t)=5400\ln(t+1),t\geq0"

"h'(t)=\dfrac{5400}{t+1}"

Since "h'(t)>0" for "t\geq0," the curve y = h(t) has no stationary points.

The height of the plane increases to infinity. as "t\to \infin."

How does the plane land?